OpenStudy (anonymous):

not good at visualizing stuff like this, any help would be wonderful!! A park in a subdivision is triangular-shaped. Two adjacent sides of the park are 533 feet and 525 feet. The angle between the sides is 53°. Find the area of the park to the nearest square foot.

OpenStudy (anonymous):

t would be wonderful if someone could help me work through it

OpenStudy (mathstudent55):



OpenStudy (anonymous):

thats cool! how did you do that?

OpenStudy (mathstudent55):

There's a drawing to help you visualize it. The adjacent sides are next ot each other and the 53 deg angle is in between them.

OpenStudy (mathstudent55):

Use draw at the bottom.

OpenStudy (anonymous):

so its just A= b x h / 2, uhm any chance you can tell me what i use for the height?

OpenStudy (mathstudent55):

The formula for the area of a triangle given two sides (a and b) and included angle C is: \[A = \frac{ 1 }{ 2 } ab \sin C\]

OpenStudy (anonymous):

shoot.. i really hate sine, cosine, and tangents :(

OpenStudy (mathstudent55):



OpenStudy (anonymous):

don't worry about it anymore it will take forever for me to understand, thanks for your help though i appreciate it alot!

OpenStudy (mathstudent55):

In drawing above, sin C = h/a (opposite over adjacent) Solve for h: h = a sin C Area of triangle = (1/2) bh, but h = a sin C, so Area = (1/2) ab sin C